# A Review of Hydrogen-Enabled Resilience Enhancement for Multi-Energy Systems

### v2

@@ Abstract
Ensuring resilience in multi-energy systems (MESs) becomes both more urgent and more challenging due to the rising occurrence and severity of extreme events (e.g., natural disasters, extreme weather, and cyber-physical attacks). Among many measures of strengthening MES resilience, the integration of hydrogen shows exceptional potential in cross-temporal flexibility, cross-spatial flexibility, cross-sector flexibility, and black start capability. Although many hydrogen-enabled MES resilience enhancement measures have been developed, the current literature lacks a systematic overview of hydrogen-enabled resilience enhancement in MESs. To fill the research gap, this paper provides a comprehensive overview of hydrogen-enabled MES resilience enhancement. First, advantages and challenges of adopting hydrogen in MES resilience enhancement are summarized. Then, we propose a resilience enhancement framework for hydrogen-enabled MESs. Under the proposed framework, existing resilience metrics and event-oriented contingency models are summarized and discussed. Furthermore, we classify hydrogen-enabled planning measures by the types of hydrogen-related facilities and provide some insights for planning problem formulation frameworks. Moreover, we categorize the hydrogen-enabled operation enhancement measures into three operation response stages: preventive, emergency, and restoration. Finally, we identify some research gaps and point out possible future directions in aspects of comprehensive resilience metric design, temporally-correlated event-targeted scenario generation, multi-type temporal-spatial cyber-physical contingency modeling under compound extreme events, multi-network multi-timescale coordinated planning and operation, low-carbon resilient planning and operation, and large language model-assisted whole-process resilience enhancement.
Index Terms:
Multi-energy systems, hydrogen-enabled resilience enhancement, resilience metrics, contingency modeling, planning, operation, extreme events, low-carbon
@@ I Introduction
Among the measures of enhancing MES resilience, the integration of hydrogen energy resources (HERs) has shown exceptional potential. To be specific, HERs refer broadly to the various technologies and systems that produce, store, transport, and utilize hydrogen as an energy carrier, including stationary HERs (e.g., electrolyzers, hydrogen tanks, hydrogen pipelines, hydrogen fuel cell-based conversion systems, and hydrogen refueling stations (HRSs)) and mobile HERs (MHERs), which include hydrogen tube trailers (HTTs)[11, 12, 13, 14, 15, 16, 17], mobile hydrogen trucks[18, 19, 20], mobile fuel cell trucks[21], hydrogen fuel cell buses (HFCBs)[22], hydrogen fuel cell vehicles (HFCVs)[23]. The above HERs can offer distinguishing features for enhancing MES resilience in aspects of energy supply and carbon emission[24, 25, 26, 27], i.e., cross-temporal flexibility, cross-spatial flexibility, cross-sector flexibility, and black start capability. Despite the promising potential of hydrogen integration, developing effective resilience enhancement strategies for hydrogen-enabled MESs (HMESs) involves significant complexity across multiple dimensions. First, defining resilience metrics for HMESs is inherently challenging due to the need to capture many aspects, e.g., load loss, carbon resilience, and physical and functional features of hydrogen systems (e.g., hydrogen availability). Second, modeling extreme event-related HESS contingencies requires careful representation of multi-stage processes (such as electrolysis, compression, storage, and reconversion) and each process is subject to distinct cyber and physical operational risks. Third, designing the schemes of hydrogen-related infrastructure planning and HMES operation faces solving challenges caused by uncertainties, mixed integer variables, multi-timescale variables, temporal-spatial coupling constraints, nonlinear and nonconvex constraints, multi-energy coupling constraints, and multi-objective function.
@@
To overcome the above challenges, researchers have defined many resilience metrics, developed some contingency models, and proposed many planning and operation methods for enhancing the resilience of HMESs[28, 29, 30, 31, 32]. However, there is a lack of a systematic overview of hydrogen-enabled resilience enhancement for MESs. To fill the research gap, we provide a comprehensive overview on leveraging hydrogen to enhance the resilience for MESs. Although there have been many review works related to power grid resilience enhancement[34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 33] with different perspectives (e.g., networked microgrids, event timeline of enhancement, cyber-physical interdependence, complex network, and cross-domain multilayer architecture), they neglect to consider the HERs. [33] is the first and only work to review the resilience of a hydrogen-powered smart grid. Although this review paper considers the hydrogen-related facilities, there are many differences between [33] and our work in aspects of objects, perspectives, contents, and research gaps. Firstly, we focus on HMESs, while [33] focuses on hydrogen-enabled smart grid. Secondly, we review the existing works from the perspective of hydrogen-enabled planning and operation. In contrast, [33] mainly focuses on the event timeline of measures, i.e., pre-event, during-event, and post-event. Thirdly, this paper summarizes advantages and challenges of adopting hydrogen in MES resilience enhancement, resilience metrics of HMES and event-oriented contingency models, which are not mentioned in [33]. Fourthly, we classify the existing resilience enhancement measures with the consideration of hydrogen pipeline hardening, more types of HERs, and multi-stage operation coordination, whereas [33] considers measures only within independent stages. Finally, we identify six research gaps and point out future directions, which are neglected in [33]. To the best of our knowledge, our work is the first survey dedicated to hydrogen-enabled resilience enhancement for MESs.
The main contributions of this paper can be summarized below.
@@
    We summarize the advantages and challenges of adopting hydrogen in MES resilience enhancement. Specifically, distinguishing features of hydrogen for resilience enhancement, resilience potential and security issues of hydrogen-related technologies, and optimization challenges of hydrogen-enabled planning and operation enhancement are summarized.
@@
    We classify the planning enhancement of HMES resilience from two aspects, i.e., single-type hydrogen facility and multi-type hydrogen facility. The considered hydrogen-related facilities consist of HESSs, hydrogen pipelines, HRSs, HFCBs, fuel cell-driven combined cooling, heat, and power (CCHP), and hydrogen-related production facilities. Moreover, the insights in aspects of uncertainty representation and formulation framework for HMES planning are summarized.
@@
    We identify six research gaps and point out future directions in aspects of comprehensive resilience metric design, temporally-correlated event-targeted scenario generation, multi-type temporal-spatial cyber-physical contingency modeling under compound extreme events, multi-network multi-timescale coordinated planning and operation, low-carbon resilient planning and operation, and large language model-assisted whole-process resilience enhancement.
@@
II Advantages and Challenges of Adopting Hydrogen in MES Resilience Enhancement
In this section, we summarize the advantages and challenges of adopting hydrogen in MES resilience enhancement. To be specific, the distinguishing features of hydrogen for MES resilience enhancement are firstly provided. Then, from the perspective hydrogen supply chain, resilience potential and security issues related to hydrogen-related technologies are summarized. Finally, the optimization advantages in planning enhancement and operation enhancement of MES resilience are analyzed.
Figure 1: Distinguishing features of hydrogen in MES resilience enhancement
@@ II-A1 Cross-temporal Flexibility
Hydrogen can provide cross-temporal flexibility[46, 47, 48]. Unlike traditional battery energy storage systems, hydrogen can be stored at a much larger scale and over longer durations, e.g., underground salt-cavern hydrogen storage can be used to provide cross-seasonal low-cost and large-scale hydrogen storage[49]. The above feature can improve MES energy resilience by providing energy backup in the case of long-duration power outages. For example, electricity and heat/cold demand in the summer and winter are typically higher than those in spring and autumn. To prepare for possible prolonged outages (e.g., lasting for more than 48 hours[50, 51, 52]) under extreme events (e.g., heat waves or ice disaster), proactive hydrogen production and storage could be conducted in spring and autumn. In addition, when produced from renewable energy, hydrogen improves the consumption rate of renewable energy. By converting excess electricity into storable hydrogen using seasonal energy storages (SHSs) and conventional HESSs, MESs can buffer supply-demand mismatches more effectively and operate with greater carbon resilience (i.e., an energy system’s ability to maintain or recover to its current and predicted emission level during and after shocks such as renewable intermittency, public health emergencies, or climate-related disruptions)[53]. As shown in [54], 100% self-sufficiency and zero carbon emission can be simultaneously achieved in distant oceanic islands with abundant renewable energy resources by coordinating SHSs and HESSs.
@@ II-A2 Cross-spatial Flexibility
Hydrogen can provide cross-spatial flexibility[46, 47, 48]. For example, hydrogen can be flexibly transported across regions through hydrogen pipelines, HTTs, or liquid hydrogen trucks, making it possible to allocate energy geographically according to regional demand and supply conditions. This spatial transfer capability becomes especially valuable in resilience enhancement when local renewable resources are insufficient or disrupted. For instance, during an earthquake, hydrogen can be transported from unaffected industrial production bases to disaster-hit areas to power temporary shelters via fuel cells, ensuring heating and lighting during prolonged grid outages. Similarly, mobile hydrogen refueling trucks can be deployed to support remote or damaged areas after extreme weather events, such as floods, by providing temporary hydrogen fuel for microgrids and emergency vehicles. By bridging spatial gaps between supply and demand, hydrogen transportation infrastructure allows MESs to dynamically reallocate resources, maintain critical services, and prevent cascading failures in interconnected energy networks, resulting in an improved energy resilience.
@@ II-A3 Cross-sector Flexibility
Hydrogen plays a crucial role in enhancing the interoperability of energy systems by acting as a versatile bridge between multiple energy carriers, including electricity, heat, gas, and transport. For example, the power-to-hydrogen technologies enable surplus electricity from renewable sources like wind and solar to be converted into hydrogen, which can then be stored and later used for electricity and heat generation or as a fuel for HFCVs in transportation sector. In addition, hydrogen can be converted to synthetic methane or blended with natural gas for transportation[31],[55],[56], which can reduce dependence on pure natural gas. The above interoperability not only enhances system efficiency, but also supports more robust and adaptive energy supply chains. Based on dynamic reconfiguration, hydrogen can facilitate resource sharing across different energy vectors, especially during extreme events. In such cases, if one subsystem experiences a failure, hydrogen can provide an alternative energy pathway, thereby ensuring the continuous operation of essential services and improving the MES resilience.
@@ II-A4 Black Start Capability
Hydrogen plays a critical role in safeguarding key infrastructures (e.g., hospitals, emergency shelters, and communication networks) during extreme events. In the event of prolonged power outages or natural gas supply failures, hydrogen-based infrastructure (such as proton exchange membrane fuel cells and hydrogen tanks) can operate off-grid and provide autonomous “black start” capabilities[57], significantly enhancing the system’s ability to self-recover without external intervention during extreme events. According to [57], black start should be implemented before network reconfiguration and load restoration could be done. Compared with black-start diesel generators, hydrogen fuel cells produce zero local emissions, operate quietly, and offer higher reliability with fewer maintenance requirements.
@@
4 Black start capability Off-grid operation of hydrogen fuel cells and tanks to restart or sustain critical infrastructure during outages Ensures rapid recovery and maintains essential services in blackout scenarios limited enhancement
@@ II-B Resilience Potential of Hydrogen-related Technologies
In the hydrogen supply chain, there are many hydrogen production pathways (e.g., electrolyzer, or steam methane reforming [58]), hydrogen storage manners (e.g., salt-cavern hydrogen storage, hydrogen tank), transportation methods (e.g., pipelines, trucks, or tube trailers), and hydrogen-to-power technologies (e.g., proton exchange membrane fuel cells, solid oxide fuel cells, and turbines), which offer different resilience capabilities in aspects of black start capability, off-grid operability, and response time, as summarized in Table III. By coordinating different hydrogen-related technologies, synergies can be achieved. For example, by combining proton exchange membrane fuel cell and on-site hydrogen tanks, fast, clean, and autonomous black start could be implemented for key infrastructures (e.g., hospitals). electrolyzers, salt cavern storage, pipelines, and hydrogen turbines, large-scale grid restoration could be implemented. Thus, a careful selection of hydrogen-related technologies tailored to the specific application scenarios is required for designing hydrogen-enabled MESs.
@@
TABLE III: Resilience Potential of Hydrogen-related Technologies
Category Technology Black start capability Off-grid operability Response time Resilience Potential Summary
Hydrogen production Electrolyzer No Yes (if coupled with renewables) Moderate (minutes) Enables distributed hydrogen production; useful in islanded scenarios but cannot self-start
Hydrogen production Steam methane reforming No No Slow (hours) Suitable for centralized supply with high continuity, but lacks flexibility and emergency use
Hydrogen storage Salt cavern storage No No Moderate (infrastructure-dependent) Suitable for seasonal, large-scale storage; requires energy infrastructure for operation
Hydrogen storage High-pressure tank No Yes Fast (minutes) Modular and deployable; enables local backup supply and fast dispatch during disruptions
Hydrogen transport Hydrogen pipeline No No Fast (seconds-minutes) Enables continuous hydrogen supply, supports regional resilience, but lacks autonomous function
Hydrogen transport HTT No Yes Moderate (minutes-hours) Highly mobile for backup fueling in off-grid or emergency zones
@@
Hydrogen utilization Proton exchange membrane fuel cell Yes Yes Fast (seconds-minutes) Ideal for critical load support, mobile/emergency backup, and black start
@@
Hydrogen utilization Solid oxide fuel cell No (needs heat-up) Yes Slow (1-2 hours) Efficient and stable for long-term off-grid power but slow to start
Hydrogen utilization Hydrogen turbines Yes (some models) Yes (with fuel supply) Moderate (minutes) Suitable for grid-scale backup and restoration, with moderate black start capability
II-C Security Issues of Hydrogen-related Technologies

### v3

@@ Abstract
Ensuring resilience in multi-energy systems (MESs) has become increasingly urgent and challenging due to the growing frequency and severity of extreme events, such as natural disasters, extreme weather, and cyber-physical attacks. Among the various approaches to enhancing MES resilience, hydrogen integration offers significant potential in cross-temporal, cross-spatial, and cross-sector flexibility, as well as black-start capability. Although considerable efforts have been devoted to this area, a systematic review of resilience enhancement in hydrogen-enabled MESs is still lacking. To address this gap, this paper presents a comprehensive review of hydrogen-enabled MES resilience enhancement. First, advantages, vulnerabilities, and challenges related to hydrogen-enabled MES resilience enhancement are summarized. Next, a resilience enhancement framework for hydrogen-enabled MESs is proposed, based on which existing resilience metrics and event-oriented contingency models are reviewed and discussed. Planning measures are then classified according to the types of hydrogen-related facilities, together with uncertainty handling methods, scenario generation methods, and planning problem formulation frameworks. In addition, operational enhancement measures are categorized into three response stages: prevention, emergency response, and restoration. Finally, research gaps are identified and future directions are discussed, including comprehensive resilience metric design, advanced extreme-event scenario generation, spatiotemporal cyber-physical contingency modeling under compound extreme events, coordinated planning and operation across multiple networks and timescales, low-carbon resilient planning and operation, and large language model-assisted whole-process resilience enhancement.
@@ I Introduction
Among the measures of enhancing MES resilience, the integration of hydrogen energy resources (HERs) has shown exceptional potential. To be specific, HERs refer broadly to the various technologies and systems that produce, store, transport, and utilize hydrogen as an energy carrier, including stationary HERs (e.g., electrolyzers, hydrogen tanks, hydrogen pipelines, hydrogen fuel cell-based conversion systems, and hydrogen refueling stations (HRSs)) and mobile HERs (MHERs), which include hydrogen tube trailers (HTTs)[11, 12, 13, 14, 15, 16, 17], mobile hydrogen trucks[18, 19, 20], mobile fuel cell trucks[21], hydrogen fuel cell buses (HFCBs)[22], hydrogen fuel cell vehicles (HFCVs)[23]. The above HERs can offer distinguishing features for enhancing MES resilience in aspects of energy supply and carbon emission as shown in Fig. 1[24, 25, 26, 27], i.e., cross-temporal flexibility, cross-spatial flexibility, cross-sector flexibility, and black start capability. Despite the promising potential of hydrogen integration, developing effective resilience enhancement strategies for hydrogen-enabled MESs (HMESs) involves significant complexity across multiple dimensions. First, defining resilience metrics for HMESs is inherently challenging due to the need to capture many aspects, e.g., load loss, carbon resilience, and physical and functional features of hydrogen systems (e.g., hydrogen availability). Second, modeling extreme event-related HESS contingencies requires careful representation of multi-stage processes (such as electrolysis, compression, storage, and reconversion) and each process is subject to distinct cyber and physical operational risks. Third, designing the schemes of hydrogen-related infrastructure planning and HMES operation faces solving challenges caused by uncertainties, mixed integer variables, multi-timescale variables, temporal-spatial coupling constraints, nonlinear and nonconvex constraints, multi-energy coupling constraints, and multi-objective functions.
@@
To overcome the above challenges, researchers have defined many resilience metrics, developed some contingency models, and proposed many planning and operation methods for enhancing the resilience of HMESs[28, 29, 30, 31, 32]. However, there is a lack of a systematic overview of hydrogen-enabled resilience enhancement for MESs. To fill the research gap, we provide a comprehensive overview of leveraging hydrogen to enhance the resilience of MESs. Although there have been many review works related to power grid resilience enhancement[34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 33] with different perspectives (e.g., networked microgrids, event timeline of enhancement, cyber-physical interdependence, complex network, and cross-domain multilayer architecture), they neglect to consider the HERs. [33] is the first and only work to review the resilience of a hydrogen-powered smart grid. Although this review paper considers the hydrogen-related facilities, there are many differences between [33] and our work in aspects of objects, perspectives, contents, and research gaps. Firstly, we focus on HMESs, while [33] focuses on hydrogen-enabled smart grid. Secondly, we review the existing works from the perspective of hydrogen-enabled planning and operation. In contrast, [33] mainly focuses on the event timeline of measures, i.e., pre-event, during-event, and post-event. Thirdly, this paper summarizes advantages and challenges of adopting hydrogen in MES resilience enhancement, resilience metrics of HMES and event-oriented contingency models, which are not mentioned in [33]. Fourthly, we classify the existing resilience enhancement measures with the consideration of hydrogen pipeline hardening, more types of HERs, and multi-stage operation coordination, whereas [33] considers measures only within independent stages. Finally, we identify six research gaps and point out future directions, which are neglected in [33]. To the best of our knowledge, our work is the first survey dedicated to hydrogen-enabled resilience enhancement for MESs.
Figure 1: Distinguishing features of hydrogen for enhancing MES resilience. Note that cross-temporal flexibility refers to the ability to store hydrogen over long durations and shift energy use from periods of surplus to periods of shortage; cross-spatial flexibility refers to the capability to transport hydrogen across regions to balance spatially uneven supply and demand; cross-sector flexibility refers to the interoperability among the electricity, heat, gas, and transport sectors through hydrogen conversion and blending; and black-start capability refers to the ability of an HMES system to self-start and restore electricity supply to critical infrastructure (e.g., hospitals, emergency shelters, and communication networks) without relying on external grid support during extreme events.
The main contributions of this paper are summarized below.
@@
    We summarize the advantages and challenges of adopting hydrogen in MES resilience enhancement. Specifically, distinguishing features of hydrogen for resilience enhancement, resilience potential and security issues of hydrogen-related technologies, and optimization challenges of hydrogen-related planning and operation are summarized.
@@
    We classify the planning enhancement of HMES resilience from two aspects, i.e., single-type hydrogen facility and multi-type hydrogen facility. The hydrogen-related facilities considered include HESSs, hydrogen pipelines, HRSs, HFCBs, fuel cell-driven combined cooling, heat, and power (CCHP), and hydrogen-related production facilities. Moreover, we summarize the classifications of uncertainty handling methods, scenario generation methods, and problem formulation frameworks for HMES planning.
@@
    We identify six research gaps and point out future directions in aspects of comprehensive resilience metric design, advanced extreme-event scenario generation, multi-type temporal-spatial cyber-physical contingency modeling under compound extreme events, multi-network multi-timescale coordinated planning and operation, low-carbon resilient planning and operation, and large language model-assisted whole-process resilience enhancement.
@@
II Advantages, Vulnerabilities, and Challenges Related to Hydrogen-enabled MES Resilience Enhancement
In this section, the distinguishing features of adopting hydrogen for MES resilience enhancement are firstly provided. Then, from the perspective of the hydrogen supply chain, potentials and vulnerabilities related to hydrogen technologies in resilience enhancement are summarized. Finally, optimization challenges associated with resilience-oriented planning and operation of hydrogen-enabled MESs are analyzed.
@@ II-A1 Cross-temporal Flexibility
This feature refers to the ability to store hydrogen over long durations and shift energy use from periods of surplus to periods of shortage[46, 47, 48]. Unlike traditional battery energy storage systems, hydrogen can be stored at a much larger scale and over longer durations, e.g., underground salt-cavern hydrogen storage can be used to provide cross-seasonal low-cost and large-scale hydrogen storage[49]. The above feature can improve MES energy resilience by providing energy backup in the case of long-duration power outages. For example, electricity and heat/cold demand in the summer and winter are typically higher than those in spring and autumn. To prepare for possible prolonged outages (e.g., lasting for more than 48 hours[50, 51, 52]) under extreme events (e.g., heat waves or ice disaster), proactive hydrogen production and storage could be conducted in spring and autumn. In addition, when produced from renewable energy, hydrogen improves the consumption rate of renewable energy. By converting excess electricity into storable hydrogen using seasonal hydrogen storage (SHS) and conventional HESSs, MESs can buffer supply-demand mismatches more effectively and operate with greater carbon resilience (i.e., an energy system’s ability to maintain or recover to its current and predicted emission level during and after shocks such as renewable intermittency, public health emergencies, or climate-related disruptions)[53]. As shown in [54], 100% self-sufficiency and zero carbon emissions can be simultaneously achieved in distant oceanic islands with abundant renewable energy resources by coordinating SHSs and HESSs.
@@ II-A2 Cross-spatial Flexibility
This feature refers to the capability to transport hydrogen across regions to balance spatially uneven supply and demand[46, 47, 48]. For example, hydrogen can be flexibly transported across regions through hydrogen pipelines, HTTs, or liquid hydrogen trucks, making it possible to allocate energy geographically according to regional demand and supply conditions. This spatial transfer capability becomes especially valuable in resilience enhancement when local renewable resources are insufficient or disrupted. For instance, during an earthquake, hydrogen can be transported from unaffected industrial production bases to disaster-hit areas to power temporary shelters via fuel cells, ensuring heating and lighting during prolonged grid outages. Similarly, mobile hydrogen refueling trucks can be deployed to support remote or damaged areas after extreme weather events, such as floods, by providing temporary hydrogen fuel for microgrids and emergency vehicles. By bridging spatial gaps between supply and demand, hydrogen transportation infrastructure allows MESs to dynamically reallocate resources, maintain critical services, and prevent cascading failures in interconnected energy networks, resulting in improved energy resilience.
@@ II-A3 Cross-sector Flexibility
This feature refers to the interoperability among the electricity, heat, gas, and transport sectors through hydrogen conversion and blending[31]. For example, the power-to-hydrogen technologies enable surplus electricity from renewable sources like wind and solar to be converted into hydrogen, which can then be stored and later used for electricity and heat generation or as a fuel for HFCVs in the transportation sector. In addition, hydrogen can be converted to synthetic methane or blended with natural gas for transportation[31],[55],[56], which can reduce dependence on pure natural gas. The above interoperability not only enhances system efficiency, but also supports more robust and adaptive energy supply chains. Based on dynamic reconfiguration, hydrogen can facilitate resource sharing across different energy vectors, especially during extreme events. In such cases, if one subsystem experiences a failure, hydrogen can provide an alternative energy pathway, thereby ensuring the continuous operation of essential services and improving MES resilience.
@@ II-A4 Black Start Capability
This feature refers to the ability of an HMES system to self-start and restore electricity supply to critical infrastructure (e.g., hospitals, emergency shelters, and communication networks) without relying on external grid support during extreme events[57]. In the event of prolonged power outages or natural gas supply failures, hydrogen-based infrastructure (such as proton exchange membrane fuel cells and hydrogen tanks) can operate off-grid and provide autonomous “black start” capabilities[57], significantly enhancing the system’s ability to self-recover without external intervention during extreme events. According to [57], black start should be implemented before network reconfiguration and load restoration are performed. Compared with black-start diesel generators, hydrogen fuel cells produce zero local emissions, operate quietly, and offer higher reliability with fewer maintenance requirements.
@@
4 Black start capability Off-grid operation of hydrogen fuel cells and tanks to restart or sustain critical infrastructure during outages Ensures rapid recovery and maintains essential services in blackout scenarios Zero carbon emission compared with diesel generators
@@ II-B Resilience Potential of Hydrogen-related Technologies
In the hydrogen supply chain, various hydrogen production pathways (e.g., electrolyzers and steam methane reforming [58]), storage options (e.g., salt-cavern hydrogen storage and hydrogen tanks), transportation modes (e.g., pipelines, trucks, and tube trailers), and hydrogen-to-power technologies (e.g., proton exchange membrane fuel cells, solid oxide fuel cells, and hydrogen turbines) provide different resilience capabilities in terms of cross-temporal flexibility, cross-spatial flexibility, cross-sector flexibility, and black-start capability, as summarized in Table III. The basis for the qualitative ratings (i.e., high, medium, and not obvious) is explained in the last column of the table. For example, salt-cavern hydrogen storage is rated as high in cross-temporal flexibility because it offers very large-scale and long-duration storage, making it particularly effective for bridging prolonged supply-demand mismatches and supporting resilience during extended disturbances. By contrast, hydrogen tanks are rated as medium because they mainly provide short-term to medium-term on-site buffering with more limited storage scale and duration. Hydrogen pipelines are also rated as medium, as they can provide a certain degree of temporal balancing through linepack; however, this capability is generally much more limited than that of dedicated geological storage, and pipelines are primarily designed for transport rather than long-duration storage. In addition, HTT is mainly characterized by cross-spatial flexibility because its principal function is to transport hydrogen from one location to another. Its cross-temporal flexibility is rated as not obvious for two reasons. First, the transportation time associated with HTT is essentially a logistics-related delivery lag rather than intentional temporal shifting enabled by storage and deferred use. In other words, simply moving hydrogen from one location to another over several hours does not provide the same type of temporal flexibility as storing hydrogen in tanks, pipelines, or salt caverns for later use. Second, although HTT may provide limited and scenario-dependent temporal support in practice (e.g., the HTT itself can serve as a form of mobile hydrogen inventory and careful dispatch may allow some degree of deferred delivery under emergency conditions), this effect is indirect, logistics-dependent, and generally weaker than the intrinsic time-shifting capability of dedicated storage technologies. It should also be noted that additional resilience benefits can be achieved through the coordinated use of different hydrogen-related technologies. For example, combining proton exchange membrane fuel cells with on-site hydrogen tanks can enable fast, clean, and autonomous black start for critical infrastructure such as hospitals. Similarly, integrating electrolyzers, salt-cavern storage, pipelines, and hydrogen turbines can support large-scale grid restoration. Therefore, careful selection and coordination of hydrogen-related technologies according to specific application scenarios are essential for the design of hydrogen-enabled MESs.
@@
TABLE III: Resilience Potential of Hydrogen-Related Technologies in MESs (H=High, M=Medium, -=Not obvious)
Category Technology Cross-temporal flexibility Cross-spatial flexibility Cross-sector flexibility Black-start capability Resilience Potential Explanations
Hydrogen production Electrolyzer - - H - Converts surplus power to hydrogen; siting is flexible but not mobile; not a generator
Hydrogen production Steam methane reforming - - H - Centralized, utility-dependent hydrogen production; strong industrial fuel support, but weak emergency agility
Hydrogen storage Salt-cavern hydrogen storage H - - - Seasonal, very-large storage; immobile; withdrawal needs powered compressors
Hydrogen storage Hydrogen tank M - - - Hours–days buffer onsite; fixed asset; enables generators but does not generate
Hydrogen transport Hydrogen pipelines M H - - Pipeline linepack; cross-spatial transport; not a generator
Hydrogen transport HTT - H - - Mobile hydrogen supply to critical loads; not a generator
@@
Hydrogen utilization Proton exchange membrane fuel cell - - H H Combined heat and power generation; Fast black-start for critical loads
@@
Hydrogen utilization Solid oxide fuel cell - - H - Combined heat and power generation; long warm-up; poor for rapid/black-start use
Hydrogen utilization Hydrogen turbines - - H H Combined heat and power generation; configured units can black-start and support restoration
II-C Vulnerabilities of Hydrogen-related Technologies